What is the wavelength of a neutron (m=1.67
Answer #1

Given, m=1.67*10-27 kg

v=7.3*104 m/s

?=h/mv

=6.62*10-34/1.67*10-27*7.3*104

=5.42*10-12 m
Answer #2

E = hv/lambda

0.5*mv^2 = h*v/lambda

lambda = 2h/mv

lambda = 2*6.626*10^-34/1.67x10^-27*7.3x10^4

lambda = 1.087x10^-11 m